OpenStudy (anonymous):

Step by step how to solve 2√ 6 * √ 8?

OpenStudy (anonymous):

Same rule....

OpenStudy (anonymous):

\[2\sqrt{6}\times \sqrt{8}\] \[2\left( 2.449 \right)\times \left( 2.828 \right)\] (4.898) * (2.828) 13.851

OpenStudy (anonymous):

Using the rule, same as 2 sqrt 48 = 2 sqrt 16 * sqrt 3 (using rule again) = 8 sqrt 3

OpenStudy (anonymous):

how did you get the * sqrt 3 part?

OpenStudy (anonymous):

It is the rule sqrt 48 = sqrt 16 * sqrt 3

OpenStudy (anonymous):

ohh. Thank you so much!

OpenStudy (anonymous):

ur welcome.